OpenStudy (anonymous):

from 1980 to 2000 the consumer price index increased from 82.4 to 172.2. If a pound of margarine cost $0.51 in 1980 and the price of margarine increased at the same rate as the CPI from 1980 to 2000 approximately how much did a pound of margarine cost in 2000?

OpenStudy (anonymous):

.51/82.4 = x/172.2 solve for x it is just a proportion (price of good in year 1)/(CPI during year 1) = (price of good in year 2)/(CPI during year 2)